# A causal random synthetic turbulent velocity field: I. The underlying Gaussian structure

### v1

@@ 4.1 Remarks concerning the Discrete Fourier Transform (DFT)
In all simulations, we consider a finite number of Fourier modes, and thus, we need to truncate in an appropriate manner the Fourier series (47) and perform an associated Discrete Fourier Transform (DFT). For full benefit of the Fast Fourier Transform (FFT) algorithm, we will be considering an even number of modes $N_{x}$ in each direction, which will furthermore be taken as a power of 2, i.e. $N_{x}=2^{p}$, with $p$ being a given integer. In this spirit, over the grid $\mathbf{x}_{\mathbf{n}}=\mathbf{n}L_{\text{\tiny{tot}}}/N_{x}$ with $\llbracket\ -N_{x}/2\ ;\ N_{x}/2-1\ \rrbracket^{3}$, here and further, we define the DFT as, for $\mathbf{k}_{\mathbf{m}}=\mathbf{m}/L_{\text{\tiny{tot}}}$ with $\mathbf{m}\in\llbracket\ -N_{x}/2\ ;\ N_{x}/2-1\ \rrbracket^{3}$,
@@ 4.2.1 Numerical scheme for the Markovian evolution ($N=1$)
Consider first the Ornstein-Uhlenbeck process given by (74), and recall that in the Space Sci. Rev. its solution is given by (79). The derivation of a numerical scheme is based on a generalization of the expression (75) of the solution $\widehat{u}^{(1),\nu}_{i}(t,\mathbf{k})$ at time $t$, depending on its initial value $\widehat{u}^{(1),\nu}_{i}(0,\mathbf{k})$ at time 0: for any nonnegative integer $m$, the solution at time $t_{m+1}=t_{m}+\delta t$ is expressed depending on the solution at time $t_{m}$ as
@@
and, using (71), the second-order moments
@@
For any $m$, consider a Gaussian white noise vector $\Gamma_{p,m}(\mathbf{x})$ and define the associated Fourier modes $\widehat{\Gamma}_{p,m}(\mathbf{k})$ following (66), i.e. set
@@
As a consequence of (67) and (68), one has the equality in distribution
@@
which can be simulated in practice by sampling a new Gaussian white noise vector and its Fourier modes at each iteration. This means that the numerical approximation defines a discrete time Markov process.
@@
4.2.2 Numerical scheme for the causal and differentiable-in-time process ($N\geq 2$)
The objective of this section is to propose a scheme for the process studied in 3.2.4 for arbitrary $N\geq 2$. Like in the case $N=1$ the aim is to define a sequence of random variables $\left(\widehat{u}_{i,m}^{(N),\nu}(\mathbf{k})\right)_{m\geq 0}$ such that one has the equality in distribution $\widehat{u}_{i,m}^{(N),\nu}(\mathbf{k})=\widehat{u}_{i}^{(N),\nu}(t_{m},\mathbf{k})$ for any nonnegative integer $m\geq 0$. The construction is more involved since the dynamics is only causal when $N\geq 2$.
@@ A Markovian formulation of the dynamics is obtained by considering the unknown
with entries $\widehat{\Gamma}_{p,m}^{(N)}(\mathbf{k}),\ldots,\widehat{\Gamma}_{p,m}^{(1)}(\mathbf{k})$ given as Fourier modes of independent Gaussian white noise vectors $\Gamma_{p,m}^{(N)}(\mathbf{x}),\ldots,\Gamma_{p,m}^{(1)}(\mathbf{x})$.
@@
The numerical approximation defines a discrete time Markov process, where at each iteration $N$ new Gaussian white noise vectors and their Fourier modes need to be computed.
@@ 5 Conclusions and perspectives
We have proposed a time evolving stochastic model of a Gaussian velocity field that reproduces several important aspects of fluid turbulence, up to second-order statistical quantities. The present version of the model is formulated in a Gaussian framework and is fully characterized by its spectral properties, or equivalently by its second-order structure function, in both space and time. To the best of our knowledge, this is the unique stochastic representation of fluid turbulence with a realistic statistical temporal structure, associated to a causal formulation, which encompasses the dimensional predictions of Tennekes (1975). The statistical signatures of this random field compare in a satisfactory manner with the observations made on DNS velocity fields, provided by the Johns Hopkins database (Li et al., 2008). We provide efficient numerical schemes, exact in probability distributions, that are straightforward to implement.
@@
Also, as we mentioned in Section 3.3 when we compared the temporal structure of the present modeled velocity field with DNS, the present approach misses an important ingredient of the sweeping effect, but presents excellent agreement from a statistical point of view. A recent proposition by Armstrong & Vicol (2025) seems to allow this property by introducing in their iterative construction of a velocity field the inverse of the Lagrangian flow map. It is thus tempting to apply such a procedure in the context of fractional Gaussian fields and explore whether it gives a realistic picture in time and if this remains affordable from a numerical point of view. Once again, we keep this aspect in mind for future investigations.
@@
[Acknowledgements]L.C. and I. S. are partially supported and J. D. L. is funded by the Simons Foundation Award ID: 651475. We thank Mokhtar Adda Bedia for his help regarding the expansion of Bessel functions in the vicinity of the origin. This work is also partially supported by Institut des Mathématiques pour la Planète Terre (Project: The role of turbulent fluctuations in the natural dispersion of aerosols, particles and pollutants).
@@ A.2 Computation of the equivalent of the kernel $F^{(N)}$ at small argument
To do so, we have to perform a development of the modified Bessel function of the second kind. We found the approach developed by Steinbock & Katzav (2024) very clear and convenient. Based on their equations (2) and (3), we write
@@
which depends on the modified Bessel function of the first kind $I_{\alpha}(x)$ with the following development
@@
Appendix B Relationship between the time-correlation of the three-dimensional Fourier modes and the one-dimension longitudinal Fourier modes
@@
Appendix C One-dimensional Fourier modes correlation function and its convergence onto a Gaussian.
We study here the Fourier modes correlation function of a one-dimensional model similar to the previous 3D one (93) and numerically show their convergence toward a Gaussian. The one-dimensional $N$-layers model writes:
$\frac{{\mathrm{d}}\widehat{u}^{(N),\nu}_{1D}(t,k)}{{\mathrm{d}}t}$ $=-\frac{\sqrt{4N}}{T_{k}}\widehat{u}^{(N),\nu}_{1D}(t,k)+\widehat{f}^{(N-1)}_{1D}(t,k)\;$ (148)
$\frac{{\mathrm{d}}\widehat{f}^{(N-1)}_{1D}(t,k)}{{\mathrm{d}}t}$ $=-\frac{\sqrt{4N}}{T_{k}}\widehat{f}^{(N-1)}_{1D}(t,k)+\widehat{f}^{(N-2)}_{1D}(t,k)\;$
$\vdots$
$\frac{{\mathrm{d}}\widehat{f}^{(2)}_{1D}(t,k)}{{\mathrm{d}}t}$ $=-\frac{\sqrt{4N}}{T_{k}}\widehat{f}^{(2)}_{1D}(t,k)+\widehat{f}^{(1)}_{1D}(t,k)\;$
${\mathrm{d}}\widehat{f}^{(1)}_{1D}(t,k)$ $=-\frac{\sqrt{4N}}{T_{k}}\widehat{f}^{(1)}_{1D}(t,k){\mathrm{d}}t+\sqrt{q^{(N)}E_{\nu}^{\text{\tiny{E,long}}}(k)}\widehat{{\mathrm{d}}W}(t,k)\;,$
where $E_{\nu}^{\text{\tiny{E,long}}}(k)$ is the longitudinal PSD as defined in (10), $T_{k}$ the characteristic time of the Fourier mode $k$ defined in (56) and $q^{(N)}$ a normalization factor defined in (94). Similarly to (95), the one-dimensional Fourier modes correlation function write in the Space Sci. Rev.:
@@
Figure 4(a) shows both numerical estimation (dots) and theoretical prediction (solid line) of Fourier modes correlation function for $N=8$ layers, in the uni-dimensional case (148) and for log-spaced Fourier modes. The inset shows the exact same curves as a function of $\tau/T_{k}$ and we observe that all curves collapse on a single nearly Gaussian function. In Figure 4(b), we show the convergence of a Fourier mode correlation function toward a Gaussian while increasing the number of layers $N$. We can observe that there is a slight but sufficiently small difference between $N=8$ and $N\rightarrow+\infty$ to consider a $8$-layers process to be numerically smooth and whose Fourier mode correlation function are almost Gaussian. The model is therefore fully able to depict a Gaussian-like Fourier mode correlation function. To lead those numerical simulations, we have chosen the same set of parameters as for the $3D$ simulations, namely $D_{3}=3.62$, $D_{2}=0.021$, $H=1/3$, $\beta=1/2$, $L=L_{\text{\tiny{tot}}}=2\pi$, $\eta_{d}=0.085$, $N_{x}=1024$, $N_{t}=5028$ and $\delta t=0.002$. The integral time scale $T_{E}$ has been defined arbitrarily by $T_{E}=\frac{L}{\sigma}$, where $\sigma^{2}$ is the variance of the process $\sigma^{2}=2\int\limits_{0}^{+\infty}E_{\nu}^{\text{\tiny{E,long}}}(k)\ {\mathrm{d}}k$. An ensemble average of the process has been made over 100 realizations to estimate numerically the Fourier modes correlation functions  in Fig. 4.

### v2

@@ 4.1 Remarks concerning the Discrete Fourier Transform (DFT)
In all simulations, we consider a finite number of Fourier modes, and thus, we need to truncate in an appropriate manner the Fourier series and perform an associated Discrete Fourier Transform (DFT). For full benefit of the Fast Fourier Transform (FFT) algorithm, we will be considering an even number of modes $N_{x}$ in each direction, which will furthermore be taken as a power of 2, i.e. $N_{x}=2^{p}$, with $p$ being a given integer. In this spirit, over the grid $\mathbf{x}_{\mathbf{n}}=\mathbf{n}L_{\text{\tiny{tot}}}/N_{x}$ with $\llbracket\ -N_{x}/2\ ;\ N_{x}/2-1\ \rrbracket^{3}$, here and further, we define the DFT as, for $\mathbf{k}_{\mathbf{m}}=\mathbf{m}/L_{\text{\tiny{tot}}}$ with $\mathbf{m}\in\llbracket\ -N_{x}/2\ ;\ N_{x}/2-1\ \rrbracket^{3}$,
@@ 4.2.1 Numerical scheme for the Markovian evolution ($N=1$)
Consider first the Ornstein-Uhlenbeck process given by (65), and recall that in the Space Sci. Rev. its solution is given by (66). The numerical scheme we use is derived from the solution (66) by computing analytically the time-evolution of the field $\widehat{u}^{(1),\nu}_{i}(t,\mathbf{k})$ at time $t_{m+1}:=t_{m}+\delta t$ as the function of the solution at time $t_{m}$:
@@
and, using (61) and the hermitian symmetry, the second-order moments
@@
For any $m$, consider a Gaussian white noise vector field $\Gamma_{p,m}(\mathbf{x})$ and define the associated Fourier modes $\widehat{\Gamma}_{p,m}(\mathbf{k})$ on a periodic domain:
@@
As a consequence of (61), one has the equality in distribution
@@
which can be simulated in practice by sampling a new Gaussian white noise vector field and its Fourier modes at each iteration. This means that the numerical approximation defines a discrete time Markov process.
@@
4.2.2 Numerical scheme for the differentiable-in-time process ($N\geq 2$)
The objective of this section is to propose a scheme for the process studied in 3.2.3 for arbitrary $N\geq 2$. Like in the case $N=1$ the aim is to define a sequence of random variables $\left(\widehat{u}_{i,m}^{(N),\nu}(\mathbf{k})\right)_{m\geq 0}$ such that one has the equality in distribution $\widehat{u}_{i,m}^{(N),\nu}(\mathbf{k})=\widehat{u}_{i}^{(N),\nu}(t_{m},\mathbf{k})$ for any nonnegative integer $m\geq 0$. When $N\geq 2$, the construction is more involved since the dynamics is Markovian only when considering the evolution of the additional underlying layers.
@@ A Markovian formulation of the dynamics is obtained by considering the unknown
with entries $\widehat{\Gamma}_{p,m}^{(N)}(\mathbf{k}),\ldots,\widehat{\Gamma}_{p,m}^{(1)}(\mathbf{k})$ given as Fourier modes of independent Gaussian white noise vector fields $\Gamma_{p,m}^{(N)}(\mathbf{x}),\ldots,\Gamma_{p,m}^{(1)}(\mathbf{x})$.
@@
The numerical approximation defines a discrete time Markov process, where at each iteration $N$ new Gaussian white noise vector fields and their Fourier modes need to be computed.
@@ 5 Conclusions and perspectives
We have proposed a time evolving stochastic model of a Gaussian velocity field that reproduces several important aspects of fluid turbulence, up to second-order statistical quantities. The present version of the model is formulated in a Gaussian framework and is fully characterized by its spectral properties, or equivalently by its second-order structure function, in both space and time. To the best of our knowledge, this is the unique stochastic representation of fluid turbulence with a realistic statistical temporal structure, associated to a Markovian formulation, which encompasses the dimensional predictions of Tennekes (1975). The statistical signatures of this random field compare in a satisfactory manner with the observations made on DNS velocity fields, provided by the Johns Hopkins database (Li et al., 2008). We provide efficient numerical schemes, exact in probability distributions, that are straightforward to implement.
@@
Also, as we mentioned in Section 3.3 when we compared the temporal structure of the present modeled velocity field with DNS, the present approach misses an important ingredient of the sweeping effect, but presents excellent agreement from a statistical point of view. A recent proposition by Armstrong & Vicol (2025) seems to allow this property by introducing in their iterative construction of a velocity field the inverse of the Lagrangian flow map. It is thus tempting to apply such a procedure in the context of fractional Gaussian fields and explore whether it gives a realistic picture in time and if this remains affordable from a numerical point of view. We believe that this realistic picture of the temporal structure would have tremendous importance when such a velocity field advects a given scalar in a passive (Falkovich et al., 2001) or active (Jossy et al., 2025) way, with direct implications on the regularity of velocity along Lagrangian trajectories (Reneuve & Chevillard, 2020). Once again, we keep these aspects in mind for future investigations.
@@
[Acknowledgements]L.C. and I. S. are partially supported and J. D. L. is funded by the Simons Foundation Award ID: 651475. L.C. is also funded by Agence National de la Recherche ANR-20-CE30-0035. We thank Mokhtar Adda Bedia for his help regarding the expansion of Bessel functions in the vicinity of the origin. This work is also partially supported by Institut des Mathématiques pour la Planète Terre (Project: The role of turbulent fluctuations in the natural dispersion of aerosols, particles and pollutants).
[…]
[Declaration of Interests]The authors report no conflict of interest.
[…]
[Note added in proof]We have become aware of the existence of a recent preprint by Awasthi et al. (2026) where is considered the one-layer velocity field $\mathbf{u}^{(1),\nu}(t,\mathbf{x})$ which Fourier mode evolution is provided in (65)
@@ A.2 Computation of the equivalent of the kernel $F^{(N)}$ at small argument
To do so, we have to perform an expansion of the modified Bessel function of the second kind. We found the approach developed by Steinbock & Katzav (2024) very clear and convenient. Based on their equations (2) and (3), we write
@@
which depends on the modified Bessel function of the first kind $I_{\alpha}(x)$ with the following expansion
@@
Appendix B Derivation of the temporal structure function and frequency spectrum in the multi-layered framework
Let us now explore the consequences of the particular temporal kernel $F^{(N)}$ given in (83) (resp. $F^{(\infty)}$ given in (84)) on the regularity in time of the corresponding velocity field $u^{(N),\nu}_{i}(t,\mathbf{x})$ (resp. $\mathbf{u}^{(\infty),\nu}_{i}(t,\mathbf{x})$). This will be illustrated with the scaling behavior of the temporal structure function, i.e. the variance of the velocity time increment (39). Using the generalized forms of $E^{\text{\tiny{E}}}(k)$ (52) and $T_{k}$, (53), the expression (2.4) reads
and when $N\to\infty$
Similarly to the temporal structure function (3.2.2) associated with the exponential kernel (68), $S_{2}^{(N),\text{\tiny{temp}}}(\tau)$ (B) and $S_{2}^{(\infty),\text{\tiny{temp}}}(\tau)$ (B) eventually undergo a transition at small scales, but for a different range of values of $H$ and $\beta$. We indeed obtain,
Similarly for $S_{2}^{(\infty),\text{\tiny{temp}}}(\tau)$, replacing $F^{(N)}$ by $F^{(\infty)}$ in (B) and computing the remaining integral using the Gaussian correlation function (84), one obtains
Note that, for a given $\beta$, the range of accessible spatial regularity $H$ is wider (i.e. $H<2\beta$) in (B) and (130) than what was found (i.e. $H<\beta$) in (70). This is a consequence of the behavior of $1-F^{(N)}(x)$ and $1-F^{(\infty)}(x)$ near the origin which go to zero as $x^{2}$, while the exponential kernel (68) predicts a much slower decay (i.e. as $x$). This is shown in (126).
For a given $\beta$, when $H$ gets large, one obtains the trivial scaling associated to a differentiable field, using the expansion of the function $F^{(N)}$ provided in (126):
A similar result is satisfied for $S_{2}^{(\infty),\text{\tiny{temp}}}(\tau)$ as a consequence of (B) when $N\to\infty$.
The scaling behaviors derived in (B) and (B) pertain to the inertial range, for which the viscosity has been set to 0 before considering any limiting behavior. We now explore the behavior of the structure function taking first the limit of vanishing scales $\tau\to 0$, and second the limit of vanishing viscosity $\nu\to 0$. When the viscosity $\nu$ is not taken to zero, instead of (B), we have
In the case $H<2\beta$, we also observe a transition in the power-law exponent of the temporal structure function, taking value $S_{2}^{(N),\text{\tiny{temp}}}(\tau)\propto|\tau|^{H/\beta}$ in the inertial range (B), and then $S_{2,\nu}^{(N),\text{\tiny{temp}}}(\tau)\propto|\tau|^{2}$ in the dissipative range (B). This transition takes place at a dissipative time scale $\tau_{d}$ defined as the very characteristic time scale at which the asymptotic expansions of both $S_{2}^{(N),\text{\tiny{temp}}}$ (obtained from (B) ) and $S_{2,\nu}^{(N),\text{\tiny{temp}}}$ (obtained from (B)) coincide. Then it is clear that $\tau_{d}$ is fully determined by viscosity and can be shown to be related to a certain power (depending on $\beta$) of the dissipative length scale $\eta_{d}$ defined in the longitudinal PSD (51).
$\widehat{F}^{(\infty)}(\omega)=\sqrt{\pi}e^{-\pi^{2}\omega^{2}},$ (135)
We show in Fig. 4(b) the convergence of the 1D Fourier modes to the Gaussian correlation function. Details are explained in the Appendices A and D Letting first viscosity vanish ($\nu\to 0$), and considering the large frequencies $\omega\to\infty$ (secondly), we obtain for $H<(2N-1)\beta$
It is worth noting that, when $N$ is finite, for $H>(2N-1)\beta$, one has
Equations (B) and (137) show that, when $H<2\beta$, even if we let the number of layers $N$ tends to infinity, the field is not differentiable in time, as it is expected in the limit of vanishing viscosity. This range of possible values of the parameter $H$ setting the regularity in time is only accessible when $N\geq 2$. Depending on the values of $N$ and $\beta$, the present approach allows to consider Hurst parameters $H$ bigger that unity, that gives rise to fields at least once differentiable in time, although physically not really realistic in a turbulent context. When $H$ gets really large, and larger than $(2N-1)\beta$, the equivalent provided in (B) says that the field is differentiable (N1)(N-1)-times, the derivatives of order N1N-1 having the regularity of the Brownian motion.
Appendix C Relationship between the time-correlation of the three-dimensional Fourier modes and the one-dimension longitudinal Fourier modes
@@ Appendix B Derivation of the temporal structure function and frequency spectrum in the multi-layered framework
Appendix D One-dimensional Fourier modes correlation function and its convergence onto a Gaussian.
We study here the Fourier modes correlation function of a one-dimensional model similar to the previous 3D one (80) and numerically show their convergence toward a Gaussian. We define the one-dimensional $N$-layers model exactly the same way as (80) without a Leray projector and by replacing $E_{\nu}^{\text{\tiny{E}}}(k)$ by $E_{\nu}^{\text{\tiny{E,long}}}(k)$. Its one-dimensional Fourier modes correlation function write in the Space Sci. Rev.:
@@
Figure 4(a) shows both numerical estimation (dots) and theoretical prediction (solid line) of Fourier modes correlation function for $N=8$ layers, in the uni-dimensional case and for log-spaced Fourier modes. The inset shows the exact same curves as a function of $\tau/T_{k}$ and we observe that all curves collapse on a single nearly Gaussian function. In Figure 4(b), we show the convergence of a Fourier mode correlation function toward a Gaussian while increasing the number of layers $N$. We can observe that there is a slight but sufficiently small difference between $N=8$ and $N\rightarrow+\infty$ to consider a $8$-layers process to be numerically smooth and whose Fourier mode correlation function are almost Gaussian. The model is therefore fully able to depict a Gaussian-like Fourier mode correlation function. To lead those numerical simulations, we have chosen the same set of parameters as for the $3D$ simulations, namely $D_{3}=3.62$, $D_{2}=0.021$, $H=1/3$, $\beta=1/2$, $L=L_{\text{\tiny{tot}}}=2\pi$, $\eta_{d}=0.085$, $N_{x}=1024$, $N_{t}=5028$ and $\delta t=0.002$. The integral time scale $T_{E}$ has been defined arbitrarily by $T_{E}=\frac{L}{\sigma}$, where $\sigma^{2}$ is the variance of the process $\sigma^{2}=2\int\limits_{0}^{+\infty}E_{\nu}^{\text{\tiny{E,long}}}(k)\ {\mathrm{d}}k$. An ensemble average of the process has been made over 100 realizations to estimate numerically the Fourier modes correlation functions in Fig. 4.